I can't seem to figure out what to do with this equation 5/8n = 3 3/4 How do I figure out n?
Answer by stanbon(48549)

5/8n = 3 3/4
(5/8)n = 15/4
Multiply both sides by 8/5 to get:
n=(15/4)(8/5)
n= 6/5
==============
Cheers,
Stan H.